OpenStudy (anonymous):

how do I get critical points from -πsin(πx)=0?

OpenStudy (anonymous):

critical point is where derivative is =0

OpenStudy (anonymous):

no i need something like cos(2sinx+1)=0 cosx=0, x=pi/2 or x= 3pi/2 sinx= -1/2 x=7pi/6 or x=11pi/6

OpenStudy (anonymous):

but i need that format with the other problem